Reports 1-1 of 1 Clear search Modify search
AdV-PSL (Laser Bench setup)
gosselin - 17:08 Tuesday 23 February 2021 (50870) Print this report
Comment to Increase of the ML pumping diode current (50755)

I took a look at the data of the past weeks.
It was difficult to get any conlusion from the data between the 12/02 (increase of the pumping current) and the 18/02. The weekend of the 13/02 14/02, we had some trouble with the lock of the RFC (log entry ). It was responsible for most of the unlocks of the weekend. Then there were several intervention in laser lab at the beginning of last week (plot 1).
The data from the 18/02 to the 23/02 were more interesting (plot 2).  Between the 18/02 and the 22/02, there were 10 fast unlocks. Only one unlock was due to the RFC.
I attached an example of those unlock (plot 3). Nothing noticeable, INJ_EOM_CORR seems to be the first one to have a glitch.
Nothing visible in the slow signal (ENV, HVAC, MAG...) for those unlocks.
We still have to properly check that we are in middle of a mode hop (log entry, but for now it seems that passing fron 2.0 A to 2.1 A of pumping current has deteriorated the situation ; while passing from 1.8 A to 2 A seemed to have improve it (see 

Images attached to this comment
Search Help